### v1

# BabyLM Turns 4: Call for Papers for the 2026 BabyLM Workshop https://babylm.github.io/

Abstract
BabyLM aims to dissolve the boundaries between cognitive modeling and language modeling. We call for both workshop papers and for researchers to join the 4th BabyLM competition. As in previous years, we call for participants in the data-efficient pretraining challenge in the general track. This year, we also offer a new track: Multilingual.
We also call for papers outside the competition in any relevant areas. These include training efficiency, cognitively plausible research, weak model evaluation, and more.
1 Introduction: BabyLM
The goals of BabyLM are to bring together multiple disciplines to answer an enduring question: how can a computational system learn language from limited inputs? Cognitive scientists investigate this question by trying to understand how humans learn their native language during childhood. Computer scientists tackle this question by attempting to build efficient machine-learning systems to accomplish this task. BabyLM aims to bridge these communities, encouraging the integration insights from cognitive science into the design of more sample-efficient language models, while also using advances in language modeling architectures to generate new hypotheses and experimental paradigms for cognitive science.
BabyLM acts as both a dedicated challenge and a workshop, for the fourth year (Warstadt et al., 2023; Hu et al., 2024; Charpentier et al., 2025). We call for submissions of original research papers at the intersection of cognitive science and language modeling, and highlight this year’s workshop theme: Going beyond English. Motivated by the recent release of BabyBabelLM (Jumelet et al., 2025), we encourage workshop submissions to engage with topics related to multilingual language acquisition and typological diversity.
For the BabyLM challenge, we again challenge participants to train language models on human-sized amounts of data, up to 100 million words. This year’s iteration will remain largely the same as previous editions, but with a few key differences, which we list below:
  • We are debuting a new Multilingual track, in which participants are tasked to train models on a multilingual dataset taken from BabyBabelLM (Jumelet et al., 2025). More details are provided in §4.1.
  • We continue to offer the Strict and Strict-small tracks, challenging participants to train on 100M and 10M words, respectively. The intermediate checkpointing requirement and the compute limitation that models may not conduct more than 10 epochs over their training data, both introduced in the past year, still remain. See §4.2 for details on how these requirements are instantiated for each track.
  • In order to streamline the competition structure, we are removing the Multimodal and Interaction tracks as dedicated tracks in the competition. These tracks are instead subsumed into the Strict and Strict-small tracks. As a result, participants may also use paired image-text data or leverage feedback from a separate teacher model during training for Strict or Strict-small. More details can be found in §4.1.
2 Key Dates
Tentative Timeline: We will accept submissions through ACL Rolling Review (ARR) or directly through OpenReview. Paper submissions to the workshop can ignore competition entry deadlines. This is our anticipated timeline. Dates will be determined based on eventual ARR and conference deadlines.
  • Mid February: CfP and training data released
  • Early April: Evaluation pipeline and baselines released
  • May 25: ARR submission deadline
  • Mid July: Direct submissions deadline
  • Early August: Direct submission reviews due, ARR commitment deadline
  • Mid August: Decisions released
  • Early September: Camera ready due
  • 24–29 Oct: Workshop @ EMNLP in Budapest (exact date TBA)
3 Non-competition Workshop Submissions
3.1 Topics
The BabyLM workshop encourages interdisciplinary submissions at the interface of language modeling, cognitive science, language acquisition, and/or evaluation. To this end, we will accept papers on a variety of topics, including but not limited to the following:
  • Data-efficient architectures and training techniques.
  • Data curation for efficient training.
  • Cognitively and linguistically inspired language modeling and evaluation.
  • Small models (and scale comparisons).
  • Relevant aspects of multimodality.
  • Interaction with or feedback from teacher models during training.
  • Second language acquisition, bilingualism or multilingualism.
Workshop Theme
The workshop will have a a theme this year: Going beyond English. Previous iterations of BabyLM have primarily focused on English, and with the introduction of the new Multilingual track we aim to inspire workshop submissions to investigate BabyLM-related topics with a focus on other languages. We hope the BabyBabelLM dataset can serve as a starting point for this, but also encourage submissions to introduce new resources that will foster progress on data-efficient language modeling across diverse linguistic settings. In particular, we welcome work that expands coverage to underrepresented and low-resource languages, explores cross-lingual transfer and multilingual training strategies, and examines how typological diversity, morphology, and script variation shape learnability under BabyLM constraints. We are also interested in benchmarks, evaluation protocols, and ethical data practices that support inclusive and culturally grounded language technologies.
3.2 Paper submission
Submissions will be made through OpenReview. Submissions can be full archival papers (or non-archival upon request) and can be up to eight pages in length. Formatting requirements will follow standards for EMNLP 2026 workshops. This includes length and anonymity requirements upon submission. Reviewing will be double-blind. As in previous iterations, we do allow dual submission; however, we do not allow dual publication.
3.3 Review & Publication
Papers submitted to the BabyLM workshop will be evaluated via a double-blind peer review process on merit and relevance. For competition participants, acceptance will be based primarily on technical soundness: we plan to reject only competition submissions that make incorrect or unjustified claims, have significant technical issues, do not reveal sufficient methodological details for replication, or demonstrate only minimal time investment. Feedback will largely be directed toward improving submissions.
4 Competition Details
4.1 Track Rules
The second BabyLM Challenge includes three competition tracks: Strict, Strict-small, and Multilingual. Note that participation in one of the competition tracks is not a prerequisite for submitting to the workshop!
New track: Multilingual
We are introducing a new Multilingual track this year, in which participants have to train models on a multilingual mixture of ‘cognitively plausible’ training data. The training data and evaluation pipeline for this track come from BabyBabelLM (Jumelet et al., 2025), which contains cognitively plausible training data for 45 different languages.
For the Multilingual track, we focus on three languages: English, Dutch, and Chinese. Focusing on the full range of BabyBabelLM would require training data beyond the scope of the BabyLM Challenge, and we therefore decided to focus on a small set of languages with broad evaluation data available. However, we highly encourage people to focus on other languages for their workshop submission: the three languages of the challenge track have been selected to streamline the challenge itself. We selected English to allow for a direct comparison to the Strict and Strict-small tracks, and to study the impact that multilingual data has on model behavior. We selected Dutch and Chinese as a way to investigate questions related to cross-lingual transfer in two ways: Dutch being a ‘typologically similar’ language to English with the same script, and Chinese being ‘typologically distant’ to both Dutch and English and with a different script.
For training, we introduce the following rule: participants are allowed to create a custom data mixture of the three languages that adds up to 100M tokens in total. Word counts per language are adjusted by their Byte Premium Arnett et al. (2024), which adjusts for variation in orthographic and morphological structure by measuring the UTF-8 encoded size needed to express a fixed amount of content. [^1] Byte Premiums are scaled with respect to English (1.0); Dutch has a premium of 1.0516, Chinese has a premium of 0.9894. What this means in practice is that 100MB of English data is equivalent to 105MB of Dutch and 98.9MB of Chinese (parallel) data. This way, it is possible to optimize the training data itself, enabling experimentation with optimal data domains and language ratios.
Evaluation will be done on a mixture of tasks assessing functional and formal linguistic competence (Mahowald et al., 2024). Depending on the task, evaluation will be conducted either through fine-tuning or via zero-shot prompting. Prior evaluations have shown that in-context learning abilities are challenging to instill into LMs of BabyLM size, which remains an open question that we keep encouraging participants to explore. Specifics on the evaluation tasks will be announced when the baseline models are released.
Continuing Tracks
The key rules for Strict and Strict-small remain unchanged from past competitions (Choshen et al., 2024; Charpentier et al., 2025). We quote:
The Strict and Strict-small tracks require that submissions be trained on a corpus of 100M words or less (in Strict) and 10M words or less (in Strict-small). These tracks do not require that participants use the official BabyLM corpus, although we will still provide an updated version of this dataset for participants interested in using it. Models in these tracks will be evaluated on language-only evaluation tasks.
This year, we have merged the Multimodal and Interaction tracks into the Strict and Strict-small tracks. This is for two primary reasons: Firstly, there were insufficient submissions to warrant standalone tracks. Secondly, none of the submissions in the Interaction track outperformed models submitted to the Strict track, and none of the submissions from Multimodal outperformed our baselines. We will still allow participants to use multimodal data or interactive training procedures if they choose, but such data or learning procedures must conform to the rules of the Strict or Strict-small tracks.
The main consequence of this change is that we now allow Strict and Strict-small models to also be trained using interactions between multiple agents during training. We will distinguish between a submission model, i.e., the participants’ entry into the competition, and an external model, i.e., a secondary model used in the training pipeline of the submission model but not submitted to the competition. From a cognitive perspective, the inclusion of an external model allows participants to simulate interactions between language learners and their caregivers or peers, as language acquisition is not passive and involves active language use from the learner (Bruner, 1985; Clark, 2020). In practice, the external model could be used in the service of the submission model’s learning in many ways. For example, this could be as simple as generating synthetic data for the submission model or providing feedback to the submission model’s outputs, which could be rendered in text or be scalar rewards for use in reinforcement learning (Martins et al., 2025).
As before, we require external models to come from a predetermined list of models available on the BabyLM website. These external models themselves may be fine-tuned or trained without restriction, however. For the Strict track, the submission model must be exposed to no more than 100M word tokens (multiple exposures allowed, e.g., epochs); this word count includes text generated by external models and pre-existing corpora. Additionally, the submission model may not generate more than 100M word tokens to be used as input for the external model. If the external model provides scalar rewards for use in RL, we also limit the number of rewards it can generate to 100M. For Strict-small, each 100M word or reward limit becomes a 10M word or reward limit instead. Finally, we note that distillation is not allowed: The external model’s tokenizer, weights, hidden states, or output distribution cannot be revealed to the submission model. If participants wish to perform knowledge distillation (Hinton et al., 2015) and use the external model’s output distribution, the training word count of the external model would count towards the submission model’s limit.
4.2 Training Requirements
As in last year’s challenge, we have epoch limits and require the submission of intermediate checkpoints:
  1. 1.
    Models submitted to the leaderboard can only be exposed to a fixed amount of input.
  2. 2.
    Intermediate model checkpoints must be submitted as well to test for learning speed and model behavior dynamics.
Below, we explain these choices in more detail.
Training Duration Limitations
Similar to last year year, we require the submission of results for models trained on a fixed amount of input (counting repeated exposures), in particular after at most 100M words for the Strict-small track and after at most 1B words for all other tracks. In most cases, this will mean after 10 epochs on the standard BabyLM corpora. However, because what counts as an epoch may differ across submissions, we instead quantify training by the number of whitespace-separated input words. For the Multilingual track, this value should be computed while taking each language’s Byte Premium into account (Arnett et al., 2024). While participants are welcome to train for longer and report this in their paper, we will only include the model checkpoints that follow these limitations in the leaderboard. Note that it is also allowed to submit a model that is trained on less data: the 100M and 1B word limits are an upper bound on data exposure. For the Strict and Strict-small tracks, the number of words seen by the submission model is considered to be the sum of the number of input words and generated tokens.
Intermediate Checkpoints
As in last year’s challenge, we require the submission of intermediate model checkpoints to the HuggingFace Hub. These checkpoints will be used for the updated evaluation pipeline (§4.4), to measure aspects related to learning efficiency and language acquisition. The checkpoints we require will be at increasing intervals: every 1M words until 10M words are seen, every 10M words until 100M words are seen, and (for the tracks other than Strict-small) every 100M words until 1B words are seen. More precise details about the evaluation of these intermediate checkpoints will be announced with the release of the evaluation pipeline.
Motivation
Before last year, we provided no such requirements. One motivation for continuing this is that the training dynamics of LMs can be compared to the learning trajectories of children, which is valuable from a cognitive modeling perspective. Furthermore, one of the conclusions of the 2024 BabyLM Challenge was that more compute is correlated with higher performance. This runs counter to the goals of BabyLM in several ways: First, one goal of BabyLM is developmentally plausible training, but children do not experience repeated exposure to their input. While we allow that memories of inputs could have an impact on learning beyond the initial exposure, we judge hundreds of repeated exposures to every input to be developmentally implausible. Second, another goal of BabyLM is to democratize pretraining research, but large numbers of training epochs require greater computational resources that are not available to all participants. As a consequence, well-funded or well-equipped research groups have a significant advantage if no limitation is applied. This advantage does not disappear with this restriction, as well-funded groups may be able to afford more hyperparameter searches and prototyping, but these efforts will at least lead to training recipes that can be reproduced in future cycles by less well-funded groups.
We have not chosen to restrict the amount of compute. While such a restriction might be ideal from the perspective of democratization, it is less clear (but by some estimates unlikely) that BabyLM submissions exceed the computation available to children (Sandberg and Bostrom, 2008). Furthermore, a requirement to compute FLOPs is more technically demanding than one to count the amount of data seen, and it could deter participation with limited additional advantages.
Finally, this restriction only applies to competition entries. Workshop papers are not required to include models with 10 or fewer training epochs, though this is, of course, encouraged.
4.3 Updated Dataset
Dataset Description # Words (multimodal data) # Words (strict data) # Images
Localized Narratives Pont-Tuset et al. (2020) Image Caption 27M 0.6M
Conceptual Captions 3M Sharma et al. (2018) Image Caption 23M 2.3M
CHILDES (MacWhinney, 2000) Child-directed speech 15M 29M
British National Corpus (BNC), dialogue portion Dialogue 4M 8M
Project Gutenberg (children’s stories) (Gerlach and Font-Clos, 2018) Written English 13M 26M
OpenSubtitles (Lison and Tiedemann, 2016) Movie subtitles 10M 20M
Simple English Wikipedia Written Simple English 7M 15M
Switchboard Dialog Act Corpus (Stolcke et al., 2000) Dialogue <<1M 1M
Total 100M 100M 2.9M
Table 1: Datasets for the strict track of the BabyLM competition. We release a text-only and multimodal dataset; either is allowed, as are outside corpora, so long as the total word count of the training corpus is less than 100M. Datasets have been detoxified following the procedure of Trhlik et al. (2026)4.3).
Recent work from last year’s Interaction Track found that BabyLMs trained on the provided Strict corpus can produce toxic generations and hate speech, which was a proven challenge for sustainable multi-turn generations between a BabyLM and a Teacher model (Salhan et al., 2025). Subsequent analysis by Trhlik et al. (2026) find that the BabyLM corpus is more toxic and hateful than training corpora used in larger models such as BERT, with a comparatively higher presence of toxic and hateful sentences, even containing racist and sexualised training data, contrary to its child-aligned disposition.
To address these issues, we release a modified detoxified training dataset based on the 2024 and 2025 training dataset (Choshen et al., 2024) at this link; see Table 1 for dataset composition statistics. This includes:
  • 100M word Strict dataset
  • 10M word Strict-small dataset
  • 100M word + image Multimodal dataset.
4.4 Evaluation
As in previous years, we will distribute an open-source evaluation pipeline, building on top of our repository from the 2025 challenge. Much of the evaluation will continue to be based on zero-shot probability comparisons of two text sequences. This year’s new track, Multilingual, will follow a similar logic to the existing evaluation for Strict. Specifically, participants will be evaluating their trained models on a combination of zero-shot and finetuning-based tasks for English, Dutch and Chinese. More details about the evaluation pipeline and the set of tasks will be released subsequently. Similar to previous years, we will release a set of hidden tasks shortly before the submission deadline, to test the robustness of models on previously unseen multilingual phenomena and domains, thereby assessing generalization beyond the released evaluation tasks.
4.5 Baselines
We will release a series of baseline models. Similar to the previous year, we will release baselines based on the winning submissions from the last year. For the Strict and Strict-small tracks, this continues to be GPT-BERT (Charpentier and Samuel, 2024). As we have merged the Interaction track into the Strict and Strict-small tracks, we will also be releasing our SimPO-based (Meng et al., 2024) preference optimization baseline from last year’s Interaction track as a baseline for the Strict and Strict-small tracks. Finally, we will continue to offer GPT-2 Small (Radford et al., 2019) as a naive, purely autoregressive baseline.
For the Multilingual track, we will be adapting our GPT-BERT and GPT-2 Small baselines to the new task setting. This will be done in a naive manner. Specifically, we will construct a 100M word dataset featuring English, Dutch and Chinese by subsampling an equal number of words from the respective dataset of each language, while taking into account each language’s Byte Premium (Arnett et al., 2024). We will then perform BPE tokenization on the resulting dataset. Training for GPT-BERT and GPT-2 Small will then proceed in the same fashion as in their Strict track counterparts.
As before, these baselines are meant to encourage participants to innovate and improve beyond existing models and approaches.
4.6 Competition Submission
Competition paper submissions will be made through OpenReview. This will include links to models and predictions, as well as links to custom datasets if applicable.
Model predictions for competition submissions must also be uploaded to a Huggingface leaderboard before the deadline. While submissions to the leaderboard after the deadline are allowed and will be displayed in the leaderboard, we will not consider these entries as participating in the competition.
What you need to submit:
  • A Huggingface link where we can download the model and its intermediate checkpoints.
  • Model predictions in a format compatible with the evaluation pipeline and Huggingface leaderboard. Our evaluation pipeline will have scripts for collating predictions for this purpose.
  • A datasheet describing the composition of the custom dataset and containing a download link (if not using a BabyLM-provided corpus)
  • When using an external model in the Strict and Strict-small tracks, the fine-tuning and distillation data for the external model (if any), and any data generated by the submission or external model.
Specific details on how and where to submit these things exactly will be shared on the BabyLM Slack and Github repository [^2] https://github.com/babylm.
5 FAQs
Can I do cool idea X?
If it is interesting, innovative, or may result in important findings, we want you to try it! If you think the rules are holding you back from submitting to the competition, please reach out to the organizers. If we find that the idea does not fit the competition’s rules, the submission can still be an interesting workshop paper.
Why doesn’t BabyLM do cool idea X?
Maybe we haven’t thought about it; please reach out.
Can papers be submitted to multiple tracks?
Yes. For example, a single paper can describe models that are submitted separately to the Strict and Multilingual tracks.
Can I submit a paper about my work?
Yes, we require that all competition submissions be accompanied by a paper, which can be up to eight pages in length (though it does not need to be). Papers will be published in an archival format. All papers can describe experiments and analyses beyond the scope of the competition.
Can I submit additional evaluation metrics?
Yes, you may submit additional evaluation metrics alongside a competition model in the Strict, Strict-small, and Multilingual tracks. This type of contribution is especially encouraged for workshop submissions.
Moreover, we accept analysis and insightful findings on previous submissions or related topics and especially welcome evaluation that works well for small models but evaluates meaningful aspects. If you believe you know of an evaluation that we should use throughout the competition, please contact us.
What training regimes are permitted?
Any training objective/regime is permitted as long as the data restrictions are followed. If you use ancillary models, for example, in the case of reranking or data augmentation, the training data for these models is counted towards your 100M word budget. This applies to all tracks; so, for example, while you can use the external model to produce POS tags, you cannot use an off-the-shelf POS tagger in your pipeline.
For evaluation purposes, we require that the model provides a function to score a sequence of words without the need for additional fine-tuning.
Are there any limits on hyperparameters?
No. But please share at the end what you found so we can learn from your efforts.
Are there any limits on the number of epochs?
As of last year’s challenge, yes. Refer to the “Training Duration Limitation” paragraph of Section 4.2 for more details.
Can I use external tools?
Yes, but if they are learned on language, their tokens are counted towards the 100M. That means one can train on the same text, both a tokenizer, a parser, an LM, etc., or on parts of the 100M, but the sum of all text seen by all training can not surpass the amount of text allowed. This raises the question of synthetic data, which is allowed under some restrictions. You may generate the 100M tokens in any legal way you like (yes, distilling or writing your own is fair, if you figure out what text facilitates learning, it is interesting regardless of how to gather such text), you may also train eventually on more than 100M words by augmentation, however, that only works in a closed system, i.e., the augmenters’ training data counts toward the limit, so, for example, training two LMs on half of the words, and then having them generate more words and training a model on both the original data and the new one is legit (and it was not tested in the previous competition, so even the example itself is interesting).
I have different modalities that can help
If it is not linguistic data, prove it, past years’ submissions did not gain from non-linguistic grounding, but we encourage such scientific questions. If it is linguistic in nature (e.g., audio), then the words should still count towards the overall number of learned words.
6 Organizing Committee
(Alphabetical by last name) Leshem Choshen, Ryan Cotterell, Mustafa Omer Gul, Jaap Jumelet, Tal Linzen, Aaron Mueller, Suchir Salhan, Raj Sanjay Shah, Alex Warstadt, and Ethan Wilcox. Feel free to contact members of the organizing committee at: leshem.choshen@mail.huji.ac.il, amueller@bu.edu, alexwarstadt@gmail.com